OpenStudy (anonymous):

does anyone know how to solve this?? 2log(base2) (x+1) = log(base2) 2+1 please help

OpenStudy (gorv):

\[2* \log_{2} (x+1)=\log_{2} 2 +1\]

OpenStudy (gorv):

so by log base property of if base and number are same thn result will be same

OpenStudy (gorv):

\[\log_{a} a=1\] alwys

OpenStudy (gorv):

\[\log_{2} 2=?????\]

OpenStudy (gorv):

@AdrianaPerez

OpenStudy (anonymous):

im a little confused, i understand that when they are both the same its 1

OpenStudy (anonymous):

What would be the steps?

OpenStudy (gorv):

\[\log_{2} 2=1\]

OpenStudy (gorv):

\[2*\log_2(x+1)=1+1=2\]

OpenStudy (gorv):

divide both side by 2

OpenStudy (gorv):

\[\log_2(x+1)=1\]

OpenStudy (gorv):

it will be 1 when both base and number are equal as base=2 so number must be =2 x+1=2 subtract 1 from both side